Grade 11Mechanics

please answer this question as soon as possible. it will help me a lot.
The velcity v of a particle moving in a straight line depends on position x as v^2=k – x^2 , where k is constant. Show that the motion of the particle is bounded within a certain region of the straight line , and find out the length of that region.

shyam
7 Years agoGrade 11

2 Answers

harishni
7 Years ago
haha
Khimraj
7 Years ago
 
the equation is
v2 + x2 = k
which is a equation of circle of radius \sqrt{k} and centre at origin.
so motion of the particle is bounded 
length of that region = 2\pi \sqrt{k}
Hope it clears..............................